The radius of curvature of a spherical mirror is 20 cm. What is its focal length?

Grade:12th pass

1 Answers

Pawan Prajapati
3 years ago
Radius of curvature, R = 20 cm Radius of curvature of a spherical mirror = 2 x Focal length (f) f = R/2 = 20/2 =10cm